solve this ohm’s law problem.

5 4. 5. Find R= 02 ISA 10Ω
Answer #1

If we treat parallel connections as functions and denote them with | sign, and if the series connection is denoted by the functional sign +, then equivalent resistance of the circuit is given by : (5|5) + (10|15)

so according to this, the resistance is:

So, total current, I = V/R = 10/8.5 = 1.17A

As resistance between AB = 5|5 =

So VAB = IRAB= 1.17*2.5 =2.92V

_______________________

As resistance between CD = 15|10 =

So VCD = IRCD= 1.17*6 =7.0V

_________________________

In a parallel connection flow of current happens to flow in the inverse ratios of the resistances, therefore,

____________________________

And,